I need these questions before my lab. Can you please list the answers in order according to number. Thank You!

You have a track, cart, cart masses, springs, wood blocks, balance, meterstick, and video analysis equipment.
Answer #1

1)Travelling down

Maximum length

Travelling back

2)If h be the height ,then Gravitational Potential Energy=Spring Potential Energy+Kinetic Energymgh=\frac{1}{2}m\dot{x}^2+\frac{1}{2}kx

3)v=\sqrt{\frac{2}{m}(mgh-\frac{1}{2}kx^2)}

From the equation it can be seen that v is maximum at x=0.

4)When then spring has maximum extension then kinetic energy is 0

So

mgh=\frac{1}{2}kx^2\\ \implies \sqrt{\frac{2gh}{k}}=x. This is the maximum extension.

Velocity is 0 at this time.